OpenStudy (anonymous):

An electric field measured near a point charge of 3.00 × 10^-6 C has a strength of 6.00 × 10^5 N/C. How far is the point from the source charge? 0.095 m , 0.21 m , 0.36 m , or 0.42 m ?

OpenStudy (michele_laino):

here we have to apply this formula: \[\Large E\left( r \right) = K\frac{Q}{{{r^2}}}\]

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

what happens next?

OpenStudy (michele_laino):

so we can write: \[\Large r = \sqrt {\frac{{KQ}}{E}} = \sqrt {\frac{{9 \times {{10}^9} \times 3 \times {{10}^{ - 6}}}}{{6 \times {{10}^5}}}} \]

OpenStudy (anonymous):

ok! so we get this? 0.212? so our solution is 0.21 m ?

OpenStudy (michele_laino):

yes! correct!

OpenStudy (anonymous):

yay! thank you!:)

OpenStudy (michele_laino):

:)